# HOMIE: Humanoid Loco-Manipulation with Isomorphic Exoskeleton Cockpit

### v1

@@ III-C1 Isomorphic Exoskeleton
To achieve accurate control and mapping of the upper limb joints of the humanoid robots, we employ an isomorphic exoskeleton as the teleoperation solution for controlling the robot’s upper body. Based on the morphology of the Unitree G1 and Fourier GR-1, our isomorphic exoskeleton design consists of a symmetric pair of arms, each with 7 DoF, corresponding to the 7 DoF of each arm of the robot (3 DoF for the shoulder, 1 DoF for the elbow, and 3 DoF for the wrist). Each joint of the exoskeleton is equipped with a DYNAMIXEL XL330-M288-T servo, which provides joint angle readings and adjustments with an accuracy of 0.09°, enabling precise joint angle mapping and initial calibration. The exoskeleton’s operational part is designed to match the length of the human arms. Considering the challenge of fully replicating the robot’s upper arm structure, we align the servos with the robot’s motor URDF joint coordinate system. By proportionally mapping the relative positions of the coordinate axes to the relative positions of the servos, we ensure that both the operator’s experience and the system’s isomorphic nature are preserved. Our isomorphic exoskeleton is fixed to the operator’s back using adjustable straps. This setup minimizes interfere with the operator’s basic movements while covering most of the upper-body motion range of the humanoid robot.
@@ III-C2 Motion-sensing Gloves
For fine teleoperation of the fingers, we adopt a joint-matching approach. Based on the Nepyone glove project [25], we design a low-cost motion-sensing glove that connects directly to the exoskeleton for assembly and use, providing up to 15 DoF for finger capture to control dexterous hands. Specifically, each finger is equipped with three sets of sensors, which map the pitch motion of the finger tip and finger pad, as well as the yaw motion of the finger pad. This setup is sufficient to enable the mapping of different dexterous hands for humanoid robots. We place Hall effect sensors and small neodymium magnets at each joint. When the joint rotates, the neodymium magnet rotates as well, thereby affecting the magnetic field sensed by the sensor and achieving the mapping of finger joint angles. Additionally, we design the glove’s microcontroller, sensor modules, and structural model. The microcontroller is mounted on the back of the hand and can be directly connected to the packaged sensors using terminal connectors, allowing for easy plugging and unplugging to reassign and modify the mapping relationships. Our motion-sensing gloves can be easily attached to and detached from different exoskeletons, offering high versatility.
@@ III-C3 Foot Pedal
IV Experiment
IV-A Locomotion Policy
@@ IV-A1 Ablation of training framework
In this section, we perform ablation experiments on the proposed upper-body pose curriculum, the height tracking reward, and the use of symmetry. All ablation experiments are conducted based on the methods described in Sec. III-B. For each setting, we use three random seeds to train policies for Unitree G1 and evaluate them in 1000 environments over a 20-second evaluation period with random upper-body poses sampled from Eq. 3 with $r_{a}\rightarrow 1$. Metrics for evaluation are tracking linear velocity error, tracking angular velocity error, tracking height error, symmetry loss and living time. The final performance for each setting is obtained by computing the average and standard deviation of the results across the three policies trained from three random seeds. All trainings are conducted on Nvidia RTX 4090 and simulated by Isaac Gym with 4096 parallel environments, where components unrelated to the ablation are kept unchanged, and only relevant parts are modified for training. Detailed parameters used in training and evaluation processes are listed in Appendix A. We mark the setting of our proposed method as ours in the following sections.
@@
Upper-body Pose Curriculum. We compare ours against two alternatives: w/o cur, which omits the curriculum and directly samples $a_{i}=\mathcal{U}(0,\mathcal{U}(0,1))$, and rand, which uses the same $r_{a}$ curriculum but replaces Eq. 3 with $a_{i}=\mathcal{U}(0,\mathcal{U}(0,r_{a}))$. Since all three methods adopt the same sampling strategy $a_{i}=\mathcal{U}(0,\mathcal{U}(0,1))$ as $r_{a}\rightarrow 1$, the final objective remains consistent, ensuring a fair comparison. The experimental results, shown in the first row of Fig. 5, reveal that ours outperforms both w/o cur and rand in linear velocity tracking, angular velocity tracking, and height error, with faster convergence and smaller errors. There is no significant difference between w/o cur and rand in the final results for these metrics. Given that the symmetry loss can reach values on the order of 20 without constraints, no significant difference is observed across the three methods in terms of symmetry loss. All three configurations achieve similar final living times, but ours and w/o cur converge more quickly. Rand, despite employing some curriculum adjustments, is limited by $r_{a}$, and values in the range $(r_{a},1]$ are not sampled during training, making it harder for the model to converge as $r_{a}$ increases. In contrast, both ours and w/o cur sample the full $[0,1]$ range from the beginning, enabling faster and more stable convergence. Thus, our curriculum approach leads to better performance compared to rand. Although w/o cur does not use a curriculum, allowing $a_{i}$ to continuously sample from $[0,1]$, the lack of difficulty smoothing leads to worse final tracking results, highlighting that our curriculum design offers a more effective training process.
@@ IV-B Teleoperation Hardware Performance
We list a series of hardware indicators for our teleoperation hardware system consisting of isomorphic exoskeleton arms, a pair of motion-sensing exoskeleton gloves, and a pedal in Tab. III. We detail their costs, with the primary expense attributed to the exoskeleton section. This is because we independently design and solder the control boards (PCBs) and sensor modules for the motion-sensing gloves and the pedal components. The acquisition frequency represents the update signal frequency measured between the hardware components of the teleoperation system and the host computer via a wired connection at a baud rate of 115200. Changing the baud rate can affect the acquisition frequency. The acquisition accuracy represents the range of angular change (in degrees) and the corresponding variation in acquisition readings, ranging from 0 to 4095( $2^{12}$). Since the mapping relationship for the motion-sensing gloves is not a clearly defined linear one, and the mapping angles for each finger joint vary, more detailed information can be found in the Appendix B.
@@ IV-C1 Real World
We deploy the trained policy on the Unitree G1 in the real world and teleoperate it to perform various loco-manipulation tasks using our isomorphic exoskeleton hardware system. The deployment code for G1 is derived from [52]. LABEL:fig:teaser (a) and LABEL:fig:teaser (c) demonstrate the robot’s capability to squat, pick objects from lower shelves, and place them on higher ones, as well as to grasp and transfer boxes between shelves utilizing its locomotion abilities. LABEL:fig:teaser (b) highlights the extensibility of our system, enabling two operators to control separate robots and collaboratively perform tasks, such as transferring apples. In LABEL:fig:teaser (d), the robot is controlled to push a 60 kg person sitting in a chair, who weighs roughly twice as much as the robot, demonstrating the robustness of the loco-manipulation system. LABEL:fig:teaser (e) illustrates how the robot uses its loco-manipulation abilities to open an oven by grasping the handle and moving backward simultaneously. LABEL:fig:teaser (f) shows that our teleoperation system is capable of performing dual-hand collaborative tasks, such as one hand passing an object to the other. LABEL:fig:teaser (g) demonstrates the robot’s ability to grasp objects from low ground, while LABEL:fig:teaser (h) shows the robot’s capability to lift and place heavy items, such as a bundle of flowers, into a box using both arms. LABEL:fig:teaser (i) demonstrates how the robot maintains balance with different upper-body poses. In all these tasks, each robot is controlled by a single operator, and the communication between the robot and operator is facilitated via Wi-Fi, without restricting the robot’s movement space. These tasks showcase the robustness of our loco-manipulation policy and HOMIE ’s ability to teleopeate humanoids perform a wide range of complex tasks in various environments.
@@
IV-C2 Simulation
@@ IV-D2 Training Setting
We adopt an end-to-end visuomotor control policy that takes images and robot proprioceptive signals as inputs and continuously outputs robot control actions. We employ a model named Seer [55], which features an autoregressive transformer architecture. Multi-view images are processed through a MAE-pretrained ViT encoder, and the features of robot proprioceptive states are extracted using an MLP. These features are subsequently concatenated into tokens. The information of these tokens are then integrated by a transformer encoder. The transformer encoder utilizes an autoregressive method to generate latent codes for controlling upper arm joints, dexterous hand movements, and height commands. The final control action output is generated by three distinct regression heads. The whole network are optimized using SmoothL1 loss. In real-world training scenarios, we configure the sequence length to 7, with both visual foresight and action prediction steps set to 3. We employ the MAE pre-trained ViT-B encoder, using bfloat16 configuration to speed up inference. This model is trained on eight A100 GPUs for 40 epoches, and we selected the checkpoint with the lowest average validation loss for evaluation.
@@ V Conclusion and Limitations
Limitations Although we have developed robust loco-manipulation policies, these policies still fall short of ensuring reliable traversal over diverse terrains. Additionally, the 15-DoF design of the motion-sensing gloves for the thumb does not fully align with human anatomy, resulting in less intuitive and smooth operation when controlling certain dexterous robotic hands. Furthermore, the current system lacks force feedback, which limits its effectiveness in applications requiring precise haptic interaction. Addressing these limitations will be a central focus of our future research efforts.
@@ A-A Network Architecture
We use a network architecture similar to that employed in a previous quadruped locomotion work called HIM [48]. The architecture consists of three main components: an estimator network $\mathcal{E}$, a follow-up network $\mathcal{N}$ that takes the output of $\mathcal{E}$ as part of its input, and a critic network $\mathcal{C}$. Together, $\mathcal{E}$ and $\mathcal{N}$ form the actor module, as illustrated in Fig. 2. All networks are implemented as 3-layer multilayer perceptrons (MLPs). Below, we describe their specific architectures, where $N_{joints}=N_{lower}+N_{upper}$, with $N_{lower}$ representing the number of lower-body joints and $N_{upper}$ the number of upper-body joints of the robot. Given that the dimensions of the input vectors are as follows: $Dim([C_{t},\omega_{t},g_{t}])=4+3+3=10$ and $Dim([v_{x},v_{y},\omega_{yaw}])=3$, we can determine the input dimensions for each network. The workflow is as follows: The sequence $O_{t-5:t}$ is fed into the encoder of $\mathcal{E}$. The encoder processes $O_{t}$ along with the ground truth values $([v_{x},v_{y},\omega_{yaw}])$ and outputs the encoded information $\hat{I}_{t}$. $\hat{I}_{t}$ is then passed to $\mathcal{N}$ and is also used for contrastive learning with the output $\hat{I}^{\prime}_{t}$ of the target network. Combining $\hat{I}_{t}$ and $O_{t}$, $\mathcal{N}$ produces the action $a_{t}$, whose dimension equals $N_{lower}$. The critic network $\mathcal{C}$ takes $O_{t}$ and the ground truth $([v_{x},v_{y},\omega_{yaw}])$ as its input.
@@
(0): Linear(in_features=$6\times(10+2\times N_{joints}+N_{lower})$, out_features=256, bias=True)
@@
(0): Linear(in_features=$10+2\times N_{joints}+N_{lower}$, out_features=256, bias=True)
@@
(0): Linear(in_features=$35+10+2\times N_{joints}+N_{lower}$, out_features=512, bias=True)
@@
(0): Linear(in_features=$3+10+2\times N_{joints}+N_{lower}$, out_features=512, bias=True)

### v2

@@ III-B4 Symmetry Utilization
Figure 4: Upper-body Exoskeleton. (a): The model architecture and physical demonstration of motion-sensing gloves. (b): The structural architecture of the upper-body exoskeleton system, comprising an isomorphic exoskeleton and motion-sensing gloves, with kinematic mapping methodology between the isomorphic exoskeleton and Unitree G1.
@@ III-C1 Isomorphic Exoskeleton
To achieve accurate control and mapping of the upper limb joints of the humanoid robots, we employ an isomorphic exoskeleton as the teleoperation solution for controlling the robot’s upper body. Based on the morphology of the Unitree G1 and Fourier GR-1, our isomorphic exoskeleton design consists of a symmetric pair of arms, each with 7 DoF, corresponding to the 7 DoF of each arm of the robot (3 DoF for the shoulder, 1 DoF for the elbow, and 3 DoF for the wrist), as shown in Fig. 4(b). Each joint of the exoskeleton is equipped with a DYNAMIXEL XL330-M288-T servo, which provides joint angle readings and adjustments with an accuracy of 0.09°, enabling precise joint angle mapping and initial calibration. Compared to incremental encoders, servo motors can store absolute position and retain the current position data even after power loss, thus eliminating recalibration upon restarts. The exoskeleton’s operational part is designed to match the length of the human arms. Considering the challenge of fully replicating the robot’s upper arm structure, we align the servos with the robot’s motor URDF joint coordinate system. We can obtain the offsets $o_{t}$ between the servos’ position angles $p_{t}$ and the robot’s joint angles $q_{t}$. Since Dynamixel servos can store absolute positions, $o_{t}$ remains fixed after assembly. Additionally, the servo disc has four symmetric holes, which causes $o_{t}$ to be an integer multiple of $\pi/2$. We use
$q_{t}=\pm k_{t}(p_{t}+\frac{n_{t}\pi}{2})+\tau_{t}$ (7)
to achieve kinematic equivalence and calibration, where the offset follows $o_{t}=\frac{n_{t}\pi}{2},\quad n_{t}\in\mathbb{Z}$ , with ±$\pm k_{t}$ is a coefficient that adjusts the direction and scale of the angle change, and $\tau_{t}$ is the additional joint angle compensation. We set $k_{t}=1$ and $\tau_{t}=0$, meaning no additional scaling or angle compensation is needed.
@@ III-C2 Motion-sensing Gloves
For fine teleoperation of the fingers, we adopt a joint-matching approach. Based on the Nepyone glove project [25], we design a low-cost motion-sensing glove that connects directly to the exoskeleton for assembly and use, providing up to 15 DoF for finger capture to control dexterous hands. Specifically, each finger is equipped with three sets of sensors, which map the pitch motion of the finger tip and finger pad, as well as the yaw motion of the finger pad. This setup is sufficient to enable the mapping of different dexterous hands for humanoid robots. We place Hall effect sensors and small neodymium magnets at each joint. When the joint rotates, the neodymium magnet rotates as well, thereby affecting the magnetic field sensed by the sensor and achieving the mapping of finger joint angles. Additionally, we design the glove’s microcontroller, sensor modules, and structural model. The microcontroller is mounted on the back of the hand and can be directly connected to the packaged sensors using terminal connectors, allowing for easy plugging and unplugging to reassign and modify the mapping relationships, as shown in as shown in Fig. 4(a). Our motion-sensing gloves can be easily attached to and detached from different exoskeletons, offering high versatility.
@@ III-C3 Foot Pedal
IV Experiments
IV-A Humanoid Whole-body Control
@@ IV-A1 Ablation of training framework
In this section, we perform ablation experiments on the proposed upper-body pose curriculum, the height tracking reward, and the use of symmetry. All ablation experiments are conducted based on the methods described in Sec. III-B. For each setting, we use three random seeds to train policies for Unitree G1 and evaluate them in 1000 environments over a 20-second evaluation period with random upper-body poses sampled from Eq. 3 with ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{a}}\rightarrow 1$. Metrics for evaluation are tracking linear velocity error, tracking angular velocity error, tracking height error, symmetry loss and living time. The final performance for each setting is obtained by computing the average and standard deviation of the results across the three policies trained from three random seeds. All trainings are conducted on Nvidia RTX 4090 and simulated by Isaac Gym with 4096 parallel environments, where components unrelated to the ablation are kept unchanged, and only relevant parts are modified for training. Detailed parameters used in training and evaluation processes are listed in Appendix A. We mark the setting of our proposed method as ours in the following sections.
@@
Upper-body Pose Curriculum. We compare ours against two alternatives: w/o cur, which omits the curriculum and directly samples $a_{i}=\mathcal{U}(0,\mathcal{U}(0,1))$, and rand, which uses the same ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{a}}$ curriculum but replaces Eq. 3 with $a_{i}=\mathcal{U}(0,\mathcal{U}(0,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{a}}))$. Since all three methods adopt the same sampling strategy $a_{i}=\mathcal{U}(0,\mathcal{U}(0,1))$ as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{a}}\rightarrow 1$, the final objective remains consistent, ensuring a fair comparison. The experimental results, shown in the first row of Fig. 6, reveal that ours outperforms both w/o cur and rand in linear velocity tracking, angular velocity tracking, and height error, with faster convergence and smaller errors. There is no significant difference between w/o cur and rand in the final results for these metrics. Given that the symmetry loss can reach values on the order of 20 without constraints, no significant difference is observed across the three methods in terms of symmetry loss. All three configurations achieve similar final living times, but ours and w/o cur converge more quickly. Rand, despite employing some curriculum adjustments, is limited by ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{a}}$, and values in the range $({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{a}},1]$ are not sampled during training, making it harder for the model to converge as ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\rho_{a}}$ increases. In contrast, both ours and w/o cur sample the full $[0,1]$ range from the beginning, enabling faster and more stable convergence. Thus, our curriculum approach leads to better performance compared to rand. Although w/o cur does not use a curriculum, allowing $a_{i}$ to continuously sample from $[0,1]$, the lack of difficulty smoothing leads to worse final tracking results, highlighting that our curriculum design offers a more effective training process.
@@ IV-B Teleoperation Hardware Performance
We list a series of hardware indicators for our teleoperation hardware system consisting of isomorphic exoskeleton arms, a pair of motion-sensing exoskeleton gloves, and a pedal in Tab. III. We detail their costs, with the primary expense attributed to the exoskeleton section. This is because we independently design and solder the control boards (PCBs) and sensor modules for the motion-sensing gloves and the pedal components. The acquisition frequency represents the update signal frequency measured between the hardware components of the teleoperation system and the host computer via a wired connection at a baud rate of 115200. Changing the baud rate can affect the acquisition frequency. The acquisition accuracy represents the range of angular change (in degrees) and the corresponding variation in acquisition readings, ranging from 0 to 4095( $2^{12}$). Since the mapping relationship for the motion-sensing gloves is not a clearly defined linear one, and the mapping angles for each finger joint vary, more detailed information can be found in the Appendix B-B.
@@ IV-C1 Real World
We deploy the trained policy on the Unitree G1 in the real world and teleoperate it to perform various loco-manipulation tasks using our isomorphic exoskeleton hardware system. We employ WiFi for communication between the cockpit and the robot. Since our system requires only 128 bytes(32-bit floats) per data packet, the measured communication latency under normal network conditions is 16 ms — a result considered acceptable for real-time control. Furthermore, we implement network techniques including checksum verification to guarantee data transmission integrity. The deployment code for G1 is derived from [52]. LABEL:fig:teaser (a) and LABEL:fig:teaser (c) demonstrate the robot’s capability to squat, pick objects from lower shelves, and place them on higher ones, as well as to grasp and transfer boxes between shelves utilizing its locomotion abilities. LABEL:fig:teaser (b) highlights the extensibility of our system, enabling two operators to control separate robots and collaboratively perform tasks, such as transferring apples. In LABEL:fig:teaser (d), the robot is controlled to push a 60 kg person sitting in a chair, who weighs roughly twice as much as the robot, demonstrating the robustness of the loco-manipulation system. LABEL:fig:teaser (e) illustrates how the robot uses its loco-manipulation abilities to open an oven by grasping the handle and moving backward simultaneously. LABEL:fig:teaser (f) shows that our teleoperation system is capable of performing dual-hand collaborative tasks, such as one hand passing an object to the other. LABEL:fig:teaser (g) demonstrates the robot’s ability to grasp objects from low ground, while LABEL:fig:teaser (h) shows the robot’s capability to lift and place heavy items, such as a bundle of flowers, into a box using both arms. LABEL:fig:teaser (i) demonstrates how the robot maintains balance with different upper-body poses. In all these tasks, each robot is controlled by a single operator, and the communication between the robot and operator is facilitated via Wi-Fi, without restricting the robot’s movement space. These tasks showcase the robustness of our loco-manipulation policy and HOMIE ’s ability to teleopeate humanoids perform a wide range of complex tasks in various environments.
@@
IV-C2 User Study
Figure 11: Teleoperation learning curves[53]. New users can quickly achieve expert speed on a new task with our hardware.
We recruit five testers with varying heights, weights, and genders to evaluate our hardware system for teleoperating Unitree G1 in Hand Over task. All novices don’t have any prior experience using the system or teleoperating humanoid robots. Each trial begins with the operator positioning both the robot and a tomato at identical designated starting points. Upon the ”Start” command, timing commences as the operator guides the robot to first grasp the tomato with one manipulator, then transfer it to the secondary manipulator. The trial concludes when the tomato is securely gripped by the receiving manipulator and the initial manipulator has fully retracted. The elapsed time from initiation to successful transfer completion is recorded as the ”completion time,” serving as the primary metric for evaluating operator proficiency with this system. After a brief tutorial, we record their completion times across five consecutive attempts. As shown in Fig. 11, the average novice time progressively approached expert-level performance despite significant physical differences among operators. This rapid improvement demonstrates our system’s intuitive usability. This result also shows out system’s strong adaptability to diverse body types with straps. Tester profiles with raw timing data are detailed in Appendix D-A, and the operational protocol in Appendix D-B.
IV-C3 Simulation
@@ IV-D2 Training Setting
Figure 13: Hardware Setup for Imitation Learning
We adopt an end-to-end visuomotor control policy that takes images and robot proprioceptive signals as inputs and continuously outputs robot control actions. We employ a model named Seer [56], which features an autoregressive transformer architecture. Multi-view images are processed through a MAE-pretrained ViT encoder, and the features of robot proprioceptive states are extracted using an MLP. These features are subsequently concatenated into tokens. The information of these tokens are then integrated by a transformer encoder. The transformer encoder utilizes an autoregressive method to generate latent codes for controlling upper arm joints, dexterous hand movements, and height commands. The final control action output is generated by three distinct regression heads. The whole network are optimized using SmoothL1 loss. In real-world training scenarios, we configure the sequence length to 7, with both visual foresight and action prediction steps set to 3. We employ the MAE pre-trained ViT-B encoder, using bfloat16 configuration to speed up inference. This model is trained on eight A100 GPUs for 40 epoches, and we select the checkpoint with the lowest average validation loss for evaluation.
@@ V Conclusion and Limitations
Limitations Our policies still fall short of ensuring reliable traversal over diverse terrains. Additionally, the 15-DoF design of the motion-sensing gloves for the thumb does not fully align with human anatomy, resulting in less intuitive and smooth operation when controlling certain dexterous robotic hands. The current system lacks force feedback, which limits its effectiveness in applications requiring precise haptic interaction. Furthermore, our exoskeleton doesn’t afford teleoperation for waist even though our policy can support arbitrary upper-body poses. Addressing these limitations will be a central focus of our future research efforts.
VI acknowledgement
This work is funded in part by the National Key R&D Program of China (2022ZD0160201), and Shanghai Artificial Intelligence Laboratory. We would like to thank Huayi Wang, Shi Zhang, Zichao Ye, Hanqing Wang, Zirui Wang, Tao Huang, and Junli Ren for their valuable discussions.
@@ A-A Network Architecture
We use a network architecture similar to that employed in a previous quadruped locomotion work called HIM [48]. The architecture consists of three main components: an estimator network $\mathcal{E}$, a follow-up network $\mathcal{N}$ that takes the output of $\mathcal{E}$ as part of its input, and a critic network $\mathcal{C}$. Together, $\mathcal{E}$ and $\mathcal{N}$ form the actor module, as illustrated in Fig. 2. All networks are implemented as 3-layer multilayer perceptrons (MLPs). Below, we describe their specific architectures, where $N_{joints}=N_{lower}+N_{upper}$, with $N_{lower}$ representing the number of lower-body joints and $N_{upper}$ the number of upper-body joints of the robot. Given that the dimensions of the input vectors are as follows: $Dim([C_{t},\omega_{t},g_{t}])=3+3+3=9$ and $Dim([v_{x},\omega_{yaw}])=2$, we can determine the input dimensions for each network. The workflow is as follows: The sequence $O_{t-5:t}$ is fed into the encoder of $\mathcal{E}$. The encoder processes $O_{t}$ along with the ground truth values $([v_{x},\omega_{yaw}])$ and outputs the encoded information $\hat{I}_{t}$. $\hat{I}_{t}$ is then passed to $\mathcal{N}$ and is also used for contrastive learning with the output $\hat{I}^{\prime}_{t}$ of the target network. Combining $\hat{I}_{t}$ and $O_{t}$, $\mathcal{N}$ produces the action $a_{t}$, whose dimension equals $N_{lower}$. The critic network $\mathcal{C}$ takes $O_{t}$ and the ground truth $([v_{x},\omega_{yaw}])$ as its input.
@@
(0): Linear(in_features=$6\times(9+2\times N_{joints}+N_{lower})$, out_features=256, bias=True)
@@
(0): Linear(in_features=$9+2\times N_{joints}+N_{lower}$, out_features=256, bias=True)
@@
(0): Linear(in_features=$35+9+2\times N_{joints}+N_{lower}$, out_features=512, bias=True)
@@
(0): Linear(in_features=$2+9+2\times N_{joints}+N_{lower}$, out_features=512, bias=True)